OpenStudy (anonymous):

i also do not understand 60 miles per hour to miles per second can you please explain too?

OpenStudy (jtvatsim):

To answer that, let's answer an easier question first... If you are going 60 miles per hour, how many miles per minute are you going?

OpenStudy (anonymous):

oh thanks!

OpenStudy (unklerhaukus):

\[1[\text {min}]=60[\text{sec}]\quad\implies\quad\frac1{60}\frac{[\text {min}]}{[\text{sec}]}=1\]\[1[\text {hr}]=60[\text{min}]\quad\implies\quad \frac{1}{60}\frac{[\text{hr}]}{[\text{min}]}=1\] \[60 \frac{[\text{mi}]}{[\text{hr}]}=60 \frac{[\text{mi}]}{[\text{hr}]}\times1\times1\] \[\qquad\qquad= 60 \frac{[\text{mi}]}{[\text{hr}]}\times\frac{1}{60}\frac{[\text{hr}]}{[\text{min}]}\times\frac1{60}\frac{[\text {min}]}{[\text{sec}]}\] \[\qquad\qquad= \cancel{60} \frac{[\text{mi}]}{\cancel{[\text{hr}]}}\times\cancel{\frac{1}{60}}\frac{\cancel{[\text{hr}]}}{\cancel{\text{min}]}}\times\frac1{60}\frac{\cancel{[\text {min}]}}{[\text{sec}]}\] \[\qquad\qquad=\frac1{60}\frac{[\text{mi}]}{[\text{sec}]}\]

OpenStudy (unklerhaukus):

can you understand that ?

OpenStudy (anonymous):

yes